OpenStudy (anonymous):

Evaluate \[\lim_{x \rightarrow 0^{+}} \frac{x^2sin\frac{1}{x}}{log(1+x)}\]

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^{+}} \frac{x^2sin\frac{1}{x}}{log(1+x)}\]\[=ln10\lim_{x \rightarrow 0^{+}} \frac{2xsin\frac{1}{x}+x^2cos\frac{1}{x}(-\frac{1}{x^2})}{\frac{1}{1+x}}\]\[=ln10\lim_{x \rightarrow 0^{+}} (1+x)(2xsin\frac{1}{x}-cos\frac{1}{x})\]Hmmm...

OpenStudy (anonymous):

lim x--->0 x sin(1/x)=0 and lim x--->0 ln(1+x)/x =1

OpenStudy (anonymous):

so answer will be zero.

OpenStudy (anonymous):

lim x--->0 x sin(1/x)=0 <- How to get this?

OpenStudy (anonymous):

since -1 <= sin(1/x) < = 1

OpenStudy (anonymous):

so lim x--->0 x sin(1/x)= 0

OpenStudy (anonymous):

does it make a sense?

OpenStudy (anonymous):

Not really. -1 ≤ sinx ≤ 1 That's clear, but in this case, it's sin(1/x), where x can be 0.

OpenStudy (anonymous):

x tends to 0+ so 1/x will tend to plus infinity. so argument of sin will be infinity but -1 <= sin(infinity)<=1

OpenStudy (anonymous):

So my question is that is sin (infinity) valid?

OpenStudy (anonymous):

yes, argument of sin can be infinity. domain of sin is whole real no.

OpenStudy (anonymous):

sin(infinity) will always have a value between -1 and +1. So (sin1/x−cos1/x) will always have a value between -1 and +1. Hence where is the problem you are facing

OpenStudy (anonymous):

Then it would become \[\lim_{y \rightarrow \infty} siny\]Can we really evaluate this limit?

OpenStudy (anonymous):

one minute...in ur solution ...u have applied hospitals rule ...right?

OpenStudy (anonymous):

Applied once. Yes.

OpenStudy (anonymous):

and this rule is applicable when the form is of 0/0 or infinity/infinity.

OpenStudy (anonymous):

so ur denominator is 0 but how ur numerator is zero?

OpenStudy (anonymous):

x^2... Hmm.. Perhaps that's where I got wrong..

OpenStudy (anonymous):

just i mean that u r doing the same whatever i did if u r applying hospitals rule.

OpenStudy (anonymous):

I say x^2, so directly put 0 into it to get 0/0. Of course, I remember x sin (1/x) = 0, but I don't know/forget how to get that.

OpenStudy (anonymous):

And what I've done doesn't mean that I'm correct..

OpenStudy (anonymous):

x sin(1/x) is zeo if you expand sin(1/x). http://upload.wikimedia.org/math/2/9/c/29cb648e96291f52707c7225630c1a17.png

OpenStudy (anonymous):

I'm sorry but I haven't learnt expanding sinx.

OpenStudy (anonymous):

can you graph xsin(1/x)

OpenStudy (anonymous):

http://mathworld.wolfram.com/MaclaurinSeries.html read this

hartnn (hartnn):

see if the explanation here for lim x sin(1/x) = 0 helps http://openstudy.com/users/waterineyes#/updates/5073f6e2e4b04aa3791e8596

hartnn (hartnn):

sin(1/x) always lies between -1 and 1 and so as x approaches zero, 0<=|x sin(1/x)| <= |x| but since x->0, |x sin(1/x)| must approach zero also.

OpenStudy (anonymous):

what kropot wrote is right

OpenStudy (anonymous):

Sandwich Theorem! That's something I heard from my friends after the quiz!

OpenStudy (anonymous):

Only problem is that I don't quite understand -1 </sin(1/x) </1 part..

hartnn (hartnn):

magnitude of sine function cannot go beyond 1 sin 90 = 1 (max) sin 270 = -1 (min) thats why those limits

OpenStudy (anonymous):

*White flag* I don't understand what I understand.. :S Thanks everyone!

OpenStudy (anonymous):

Is infinity included the domain of sine?

OpenStudy (anonymous):

*included in

hartnn (hartnn):

yes. why not....but u cannot determine the value of sin(infinity)

OpenStudy (anonymous):

hey, see the graph of sine function ..... its obvious

OpenStudy (kropot72):

\[\lim_{x \rightarrow 0}=\frac{x ^{2}\sin \frac{1}{x}}{\log_{} (1+x)}=\frac{\frac{x ^{2}}{x}}{x}=1\] the reasons being: (1) as x ---> 0, sin 1/x = 1/x (2) as x ---> 0, log (1 + x) = x

OpenStudy (anonymous):

u cant say lim x-->0 sin(1/x)= 1/x??

OpenStudy (anonymous):

i think u r using for small x sinx=x but how u can approximate sin(1/x) with 1/x?

OpenStudy (anonymous):

limit wont be 1, it'll be 0.

OpenStudy (anonymous):

lim x--->0 x sin(1/x)=0 and lim x--->0 ln(1+x)/x =1 so lim x-->0 x^2 sin(1/x)/ ln(1+x) =0

OpenStudy (kropot72):

@akash123 Sorry. My bad :(

OpenStudy (anonymous):

hey it's fine...:) plz no sorry.

OpenStudy (anonymous):

best way to calculate simple limits is by using graphs....

OpenStudy (anonymous):

Hmmm... Please ignore the following.. ------------------------------------------------------------ sinx: Domain: (-∞, ∞) Range: [-1, 1] So, -1 ≤ sinx ≤ 1 for (-∞, ∞) -1 ≤ sin (1/x) ≤ 1 -x^2 ≤ x^2 sin (1/x) ≤ x^2 \[\lim_{x \rightarrow 0^+}-x^2 = 0\]\[\lim_{x \rightarrow 0^+}x^2 = 0\]By Sandwich Theorem, \[\lim_{x \rightarrow 0^+}x^2 sin \frac{1}{x} = 0\] For cosine cosx: Domain: (-∞, ∞) Range: [-1, 1] So, -1 ≤ cosx ≤ 1 for (-∞, ∞) -1 ≤ cos (1/x) ≤ 1 How do you evaluate the limit of cos(1/x) when x-> 0+ ?

OpenStudy (anonymous):

lim x-->0 cos(1/x) indeterminate

OpenStudy (anonymous):

take 1/x=y then when y tends to infinity cos y tends to? you yourself know this

hartnn (hartnn):

yeah, thats inderterminate and u don't need it here

OpenStudy (anonymous):

That's the problem! \[=ln10\lim_{x \rightarrow 0^{+}} (1+x)(2xsin\frac{1}{x}-cos\frac{1}{x})\] I can't evalute the cos(1/x)

hartnn (hartnn):

\(\huge \lim_{x \rightarrow 0^{+}} \frac{xsin\frac{1}{x}}{log(1+x)/x}=\frac{0}{1}=0\)

hartnn (hartnn):

distribute the limits in num, and denom

hartnn (hartnn):

num limit through sandwich th.

OpenStudy (anonymous):

How do you get that 1?

hartnn (hartnn):

denominator lim through l'hopital or u can use it as a standard formula

hartnn (hartnn):

through L'hopital's its 1/(1+x) when x->0+ = 1

hartnn (hartnn):

num derivative = 1/1+x denom derivative = 1, then at x=0+, the limit =1

OpenStudy (anonymous):

.............. fraction in denominator...... can we still apply L'Hopital's rule for that? (Probably your answer is yes......)

hartnn (hartnn):

aaaand it is yes

hartnn (hartnn):

u just need to check for 0/0 in denominator

OpenStudy (anonymous):

(Cry.....) I am sooooooooo stupiddddd!!! Thanks everyone!!!

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0^{+}} \frac{x^2sin\frac{1}{x}}{log(1+x)}\] \[=\lim_{x \rightarrow 0^{+}} \frac{xsin\frac{1}{x}}{\frac{log(1+x)}{x}}\] \[= \frac{\lim_{x \rightarrow 0^{+}}xsin\frac{1}{x}}{\lim_{x \rightarrow 0^{+}}\frac{log(1+x)}{x}}\] Consider \(\lim_{x \rightarrow 0^{+}}xsin\frac{1}{x}\) \[\lim_{x \rightarrow 0^{+}}xsin\frac{1}{x}\] \[-1 \le sin\frac{1}{x} \le 1\] \[-x \le xsin\frac{1}{x} \le x\] \[lim_{x \rightarrow 0^+} –x = lim_{x \rightarrow 0^+} x = 0\] By Sandwich Theorem, \[lim_{x \rightarrow 0^+} xsin\frac{1}{x} = 0\] Consider \(\lim_{x \rightarrow 0^{+}}\frac{log(1+x)}{x}\) By L’Hopital’s Rule \[\lim_{x \rightarrow 0^{+}}\frac{log(1+x)}{x}\] \[=\lim_{x \rightarrow 0^{+}}\frac{\frac{1}{1+x}}{1}\] \[=\lim_{x \rightarrow 0^{+}}\frac{1}{1+x} = 1\] So, \[\frac{\lim_{x \rightarrow 0^{+}}xsin\frac{1}{x}}{\lim_{x \rightarrow 0^{+}}\frac{log(1+x)}{x}}\] \[=\frac{0}{1}=0\]

OpenStudy (anonymous):

good work...:)

OpenStudy (anonymous):

One minor question, when we do differentiation for common log, do we have to convert into natural log?

hartnn (hartnn):

in questions relating limits, integration.... log is considered as natural log (ln) unless its mentioned as log here is common log , in which case u need to convert it into natural log ln

hartnn (hartnn):

so the answer to your question is YES

OpenStudy (anonymous):

Awww... That means to be safe, I should convert it first, right? (Since usually whenever I see log, it's common log)

hartnn (hartnn):

you should convert only if it is mentioned that log is "common" log else consider it as ln as you did it in this question *only for problems on limits, diff., integration

OpenStudy (anonymous):

Okay, thanks! So, for logx, there is no difference, but for \(\log_2x\), there is a problem, right?

hartnn (hartnn):

yes, covert log2 x to ln x here

OpenStudy (anonymous):

Nice! Thanks!!!~